# Passing the Buck to AI: How Individuals’ Decision-Making Patterns Affect Reliance on AI

### v1

@@ 5.1. Implications for Designing AI Systems and Explainable AI Research
Delay AI suggestions for buckpassers to reduce tendency of deferring decision-making to AI
Our findings show that those prone to deferring decisions to others are also more likely to rely on the AI. The relatively short time they spent looking at the AI evaluations, compared to participants who scored low in buckpassing, further suggests that buckpassers may not think for themselves in these situations; the immediate access to the AI suggestions (as is most common in human-AI studies and AI decision-support systems in various contexts (Chen et al., 2023; Morrison et al., 2023; Levy et al., 2021; Yin et al., 2019; Cao et al., 2023; Lu et al., 2024)) seems to be an easy way to lower cognitive effort, which then results in an overreliance on the AI. Buçinca et al. (2021) found that cognitive forcing functions—forcing people to make their own decisions before seeing an AI’s suggestions—can effectively reduce their overreliance.Hence, delaying the option to seek AI assistance could be especially beneficial for buckpassers, encouraging them to form an opinion themselves before seeing the AI’s suggestions.
Note that in our study, the impact of decision-making patterns on the accuracy of people’s decisions was negligible. This was likely due to the relatively easy task, which meant that even when heavily relying on the AI, participants didn’t have significantly worse decision outcomes. However, the increased reliance on the AI that we saw in our buckpassing participants suggests that delaying the AI suggestions is crucial in high-stake decision-making tasks. AI designs should therefore detect who may be prone to buckpassing behavior, such as by detecting people’s decision-fatigue using their cursor entropy as suggested in (Reinhardt and Hurtienne, 2018) or the amount of time they tend to spend reading AI-generated explanations). Cognitive forcing functions could then be applied in a personalized manner and reactive to an individual’s decision-fatigue to reduce overreliance.
@@
Our study found that buckpassers tend to spend less, and vigilant people more time, when reviewing AI explanations, posing challenges to the design of AI explanations. The format of AI explanations in our study was text-based and lengthy. This format might induce further cognitive load and stress for buckpassers, resulting in their reduced time investment in AI explanations. Prior research found if decisions are made under high cognitive load, it can lead to higher overreliance on AI (Zhang et al., 2024). Thus, identifying presentation formats of AI suggestions that reduce cognitive load and decision stress may enourage buckpassers to evaluate AI explanations carefully.
Chen et al. (2023) found that for individuals who rely on their intuition in decision-making, providing example-based scenarios as AI explanations can effectively calibrate appropriate reliance on AI. Psychology research on cognitive load theory (CLT) and e-learning has identified various information presentation techniques to reduce individuals’ cognitive load (Kirschner, 2002; Van Merrienboer and Ayres, 2005; Sweller, 2011). Mousavi et al. (1995) found that presenting information in both audio and visual format leads to less cognitive load than in visual format alone. As such, designers could consider showing AI explanations to users in both visual and audio formats to lower their cognitive load. Redundant information is also shown to increase individuals’ cognitive load and experts and novices have different perceptions of redundancy (Sweller, 2011). In AI suggestions, designers should personalize information based on individuals’ domain knowledge level and avoid providing information that is commonly known among domain experts.
Provide users with diverse options for presentation type of AI suggestions
Despite the effectiveness of interventions in cultivating appropriate reliance on AI, research in explainable AI has focused on explanations for binary decision-making tasks; however, the information that conversational AI systems, such as ChatGPT, provide often use one-size-fits-all presentation format. The varied amount of time investment could result in a different level of caution when processing information from ChatGPT. While these lengthy paragraphs might suit the information-seeking habits of vigilant people, they might not be processed carefully by individuals who tend to be buckpassers. As such, developers should consider providing features that enable users to select their own preferred presentation format based on their own information-seeking habits. For example, if users may adjust the length of text and amount of details (i.e. summary of references or concrete examples) used in AI suggestions, vigilant individuals may choose to obtain more comprehensive information whereas buckpassers could request brief explanations.
@@ 6. Limitations and Future Work
We acknowledge several limitations of this study. Firstly, the observed behaviors in the design of this study may not generalize to other domains. Our study was designed to capture participants’ collaborative decision-making behavior with an AI in a task that required them to accurately identify whether a set of nutrition statements were facts or myths. Choosing nutrition as a domain has the advantage that the study was broadly accessible to diverse participants, but inaccuracies have naturally less severe consequences than tasks in, e.g., the medical domain. We additionally chose to conduct an online study with volunteer participants who were offered personalized performance feedback in exchange for study completion. This usually means participants are intrinsically motivated to do well; however, the consequences of mistaking a nutrition statement as true when it was false (and vice versa) are mild when compared to a real-world situation. Future work is needed to validate the findings with tasks that carry greater consequences for participants in other real-world scenarios.
Secondly, our study did not fully capture individuals’ various reliance outcomes. Prior studies on AI-assisted decision-making often investigate individuals’ overreliance or underreliance on AI (Cao et al., 2024; Vasconcelos et al., 2023). Our design examines individuals’ interactions with AI suggestions by focusing on whether they seek out AI suggestions and how they rely on the AI overall. Future studies may examine reliance outcomes together with the observed behaviors in our study to provide a holistic understanding of individuals’ reliance behaviors.
@@
Thirdly, our study design does not fully capture users’ holistic experiences with AI systems, such as ChatGPT, through chatbot interfaces since users did not have a choice to see more AI explanations if they were not satisfied with the output. We chose such a design to ensure AI suggestions are consistent for each participant, yet we hope future studies could explore experiment designs that closely mimic interactions with generative AI systems. Meanwhile, our design of asking users’ to indicate their adoption of AI suggestion by providing an option “Use the AI Decision” had very few engagement click despite the high reliance suggested by other measures. We suspect this occurred due to an unintuitive display of options for users to choose from or a motivation to preserve autonomy in making the final decision. In future work, we will also improve our designs of interface features that are more closely matched with users’ behaviors. It’s also worth noting that our design of the study does not provide the opportunity to search for information from the web. In real-world scenarios, users could potentially navigate other information sources for more information before they make their final decisions. Future work could further integrate this option to examine how users’ behaviors are influenced by their decision-making patterns in a more dynamic setting. Notably, while we did not examine the effect of procrastination in our study, future work that integrates more dynamic interactions shall investigate the behaviors of individuals who tend to procrastinate in their decision-making.
@@ 7. Conclusion
With the increasing use of conversational AI such as ChatGPT for information-seeking and decision-making, it is essential to understand how people vary in their interactions with AI suggestions in the decision-making process. Through an online study, we asked participants (n=810) to evaluate the factuality of nutrition-related statements with the option to seek AI suggestions (decisions and explanations). We found that people who tend to defer decisions to others (buckpassers) are more likely to seek AI suggestions yet spend less time evaluating these suggestions and reported a higher level of reliance on AI when evaluating nutrition information than those scoring low on buckpassing. In contrast, vigilant decision makers tended to more carefully scrutinize the AI’s information than those scoring low on vigilance. Drawing insight from psychology research on decision-making, our study suggests that individuals’ decision-making patterns implicate not only human-human interactions but also human-AI interactions. In particular, these findings expand the current research horizon of AI-assisted decision-making by underscoring the importance of individual cognitive processes. As more AI-driven systems are being developed and integrated into our everyday lives, these findings shed light on the importance of individual cognitive factors, providing new insights for the future development of AI technologies.

### v2

@@ 5.1. Implications for Designing AI Systems and Explainable AI Research
Account for the differential effects of cognitive forcing functions
Our findings showed that those prone to deferring decisions to others are also more likely to seek out AI suggestions. The relatively short time they spent looking at the AI explanations, compared to participants who scored low in buckpassing, further suggests that buckpassers may not engage as strongly with this information; the immediate access to the AI suggestions (as is most common in human-AI studies and AI decision-support systems in various contexts (Chen et al., 2023; Morrison et al., 2023; Levy et al., 2021; Yin et al., 2019; Cao et al., 2023; Lu et al., 2024)) seems to be an easy way to lower cognitive effort, which may lead to overreliance on the AI. Cognitive forcing functions—like presenting AI suggestions on demand or forcing people to make their own decisions before seeing an AI’s suggestions—are effective strategies researchers have identified in reducing overreliance (Buçinca et al., 2020). Yet it remains unclear how individuals differ in interacting with each kind of cognitive forcing function. Our results suggest that decision-making patterns also influence individuals’ engagement with AI suggestions when they are presented on demand. In addition, since buckpassers often defer decision-making to others due to experienced stress, implementing cognitive forcing functions may lead to more mental demand and stress in decision-making. Thus, designers should prioritize relatively low-effort cognitive forcing functions and couple them with motivational nudges (Caraban et al., 2019; Vasconcelos et al., 2023) (e.g., by reminding the potential consequences of overreliance behaviors (Bo et al., 2025)) to encourage active engagement among this group.
@@
Our study found that buckpassers tend to spend less, and vigilant people more time, when reviewing AI explanations, posing challenges to the design of AI explanations. The format of AI explanations in our study was text-based and lengthy. This format might induce further cognitive load and stress for buckpassers, resulting in their reduced time investment in AI explanations. Prior research found that if decisions are made under high cognitive load, they can lead to higher overreliance on AI (Zhang et al., 2024). Thus, similar to cognitive forcing functions, identifying presentation formats of AI suggestions that reduce cognitive load and decision stress may encourage buckpassers to evaluate AI explanations carefully. Chen et al. (2023) find that for individuals who rely on their intuition in decision-making, providing example-based scenarios as AI explanations can effectively calibrate appropriate reliance on AI. Psychology research on cognitive load theory (CLT) and e-learning has identified various information presentation techniques to reduce individuals’ cognitive load (Kirschner, 2002; Van Merrienboer and Ayres, 2005; Sweller, 2011). Mousavi et al. (1995) found that presenting information in both audio and visual format leads to less cognitive load than in visual format alone. As such, designers could consider showing AI explanations to users in both visual and audio formats to lower their cognitive load. Redundant information is also shown to increase individuals’ cognitive load (Sweller, 2011). However, experts and novices have different perceptions of redundancy; in AI suggestions, designers should personalize information based on individuals’ domain knowledge and avoid providing information that is commonly known among domain experts.
Provide users with diverse options for the presentation type of AI suggestions
Despite the effectiveness of interventions in cultivating appropriate reliance on AI, research in explainable AI has focused on explanations for binary decision-making tasks; however, AI systems often generate information in a limited presentation format and a lack of user control. The varied amount of time investment could result in a different level of caution when processing information from AI. While these lengthy paragraphs might suit the information-seeking habits of vigilant people, they might not be processed carefully by individuals who tend to be buckpassers. As such, developers should consider providing features that enable users to select their own preferred presentation format based on their own information-seeking habits. For example, if users may adjust the length of text and amount of details (i.e. summary of references or concrete examples) used in AI suggestions, vigilant individuals may choose to obtain more comprehensive information whereas buckpassers could request brief explanations.
@@ 6. Limitations and Future Work
We acknowledge several limitations of this study. First, the observed behaviors in the design of this study may not generalize to other domains and tasks. Our study was designed to capture participants’ collaborative decision-making behavior with an AI in a task that required them to accurately identify whether a set of nutrition statements are facts or myths. Choosing nutrition as a domain has the advantage that the study was broadly accessible to diverse participants, but inaccuracies have naturally less severe consequences than tasks in, e.g., the medical domain. We additionally chose to conduct an online study with volunteer participants who were offered personalized performance feedback in exchange for study completion. This usually means participants are intrinsically motivated to do well; however, the consequences of mistaking a nutrition statement as true when it was false (and vice versa) are mild when compared to a real-world situation. While similar effects of decision-making patterns may also occur in other domains and general-purpose tasks, these effects could be influenced by task differences. For tasks that are more complex, these effects may be exacerbated by the complexity of the task (Bansal et al., 2019). For tasks that have less objective ground truth, these effects may still persist yet decision outcomes may change based on individuals’ expertise. Future work is needed to confirm how task complexity and outcome severity shape these effects. We encourage researchers to adopt similar study designs in other settings and professional domains to better understand these dynamics.
Secondly, our study did not fully capture individuals’ various reliance outcomes. Prior studies on AI-assisted decision-making investigated individuals’ appropriate reliance (overreliance or underreliance) on AI (Cao et al., 2024; Vasconcelos et al., 2023). These measures were obtained in a study design that involves two-stage decision-making (e.g., individuals are asked to make decisions first before they are exposed to AI suggestions, and then make final decisions after seeing AI suggestions). Our study adopted a design setting different from two-stage decision-making, because we focused on whether participants seek out AI suggestions and how they perceive their reliance on the AI overall. This means that our study outcome is only partially comparable to prior findings on AI overreliance. Specifically, while we observed that participants who scored high in buckpassing tend to perform worse than those who scored low in buckpassing when both were exposed to incorrect AI decisions, this difference was not significant. Therefore, while participants with high buckpassing tendency were significantly more likely to seek out AI suggestions, we cannot draw firm conclusions about how this tendency influences their overall decision performance. Future work is needed to disentangle whether buckpassers tend to overrely on AI in addition to being more likely to seek AI assistance.
@@
Thirdly, our study design did not fully capture users’ holistic experiences with AI systems, such as ChatGPT, through chatbot interfaces since users did not have a choice to see more AI explanations if they were not satisfied with the output. We chose such a design to ensure AI suggestions are consistent for each participant, yet we hope future studies could explore experiment designs that closely mimic interactions with generative AI systems. Meanwhile, our design of asking users’ to indicate their adoption of AI suggestion by providing an option “Use the AI Decision” had very few engagement click despite the high reliance suggested by other measures. We suspect this occurred due to an unintuitive display of options for users to choose from or a motivation to preserve autonomy in making the final decision. In future work, we will also improve our designs of interface features that are more closely matched with users’ behaviors. It’s also worth noting that our design of the study does not provide the opportunity to search for information from the web. In real-world scenarios, users could potentially navigate other information sources for more information before they make their final decisions. Future work could further integrate this option to examine how users’ behaviors are influenced by their decision-making patterns in a more dynamic setting. Notably, while we did not examine the effect of procrastination in our study, future work that integrates more dynamic interactions shall investigate the behaviors of individuals who tend to procrastinate in their decision-making.
@@ 7. Conclusion
With the increasing use of conversational AI such as ChatGPT for information-seeking and decision-making, it is essential to understand how people vary in their interactions with AI suggestions in the decision-making process. Through an online study, we asked participants (n=810) to evaluate the factuality of nutrition-related statements with the option to seek AI suggestions (decisions and explanations). We found that people who tend to defer decisions to others (buckpassers) are more likely to seek AI suggestions yet spend less time evaluating these suggestions and reported a higher level of reliance on AI when evaluating nutrition information than those scoring low on buckpassing. In contrast, vigilant decision makers tend to more carefully scrutinize the AI’s information than those scoring low on vigilance. Drawing insight from psychology research on decision-making, our study suggests that individuals’ decision-making patterns implicate not only human-human interactions but also human-AI interactions. In particular, these findings expand the current research horizon of AI-assisted decision-making by underscoring the importance of individual cognitive processes. As more AI-driven systems are being developed and integrated into our everyday lives, these findings shed light on the importance of individual cognitive factors, providing new insights for the future development of AI technologies.
Acknowledgements.
This work is funded by the National Science Foundation under Grant No.2230466.
@@ Appendix B The Melbourne Decision-Making Questionnaire
The Melbourne Decision-Making Questionnaire by Mann et al. [44], used to assess participants’ decision-making patterns. Items are answered on a 3-point scale: ’Not true for me,’ ’Sometimes true for me,’ or ’True for me.’ - Vigilance (Items 1–6): Statements focus on considering alternatives, analyzing disadvantages, collecting information, and being deliberate and clear in decision-making. - Buckpassing (Items 7–12): Statements highlight avoiding or deferring decisions, preferring others to decide, and avoiding responsibility. - Hypervigilance (Items 13–17): Statements express feeling pressured, pessimistic, or anxious about decisions, and struggling with rushed or uncertain choices.
@@ B.1. Study Design
An image showing an example of a result page presented to participants in a study. The page provides feedback on participants’ performance in a food knowledge task, showing that they answered 5 out of 10 questions correctly. It lists the questions they got wrong and the correct answers, such as ”Diet should be adapted to a person’s blood group. The answer is: This is a myth,” and ”The day should always start with breakfast. The answer is: This is a fact.” Additionally, the page explains that ChatGPT was used to generate AI-based explanations and decisions, noting that some may be inaccurate. Specific questions where AI suggestions were inaccurate are highlighted, including topics like dietary adaptation to blood groups and the caloric comparison between normal and sweet potatoes.
@@ B.2. Factor Analysis for Decision-making Patterns
A factor loading diagram illustrating the three decision-making patterns based on the Melbourne Decision-Making Questionnaire. The diagram includes three latent factors: PA1 (Buckpassing), PA2 (Vigilance), and PA3 (Hypervigilance). Items related to buckpassing (e.g., questions-buckpassing-p1 through p6) have strong loadings on PA1, items related to vigilance (e.g., questions-vigilance-p1 through p6) load on PA2, and items related to hypervigilance (e.g., questions-hypervigilance-p1 through p5) load on PA3. The inter-factor correlation between PA1 and PA3 is depicted with a value of 0.7. Each loading value (ranging from 0.5 to 0.8) represents the strength of the relationship between individual items and their respective factors, with higher values indicating stronger associations.
@@ Appendix C Demographics Information of Participants
A histogram showing the age distribution of 810 participants. The x-axis represents age, ranging from 12 to 90 years, while the y-axis represents the count of participants. The distribution is skewed towards younger ages, with the highest count in the 15–25 age range. The number of participants decreases steadily with age, showing a sharp decline after 30 and tapering off significantly beyond 60. Only a few participants are above 70 years old.
@@ B V H A HE CE GE DKE
Figure 9. Distribution of Conditional Agreement with AI Decision Rate: conditional agreement rate is participants’ agreement rate conditioned on whether they chose to see AI decisions. There are 395 out of 810 participants (49%) who agreed with AI decisions when they chose to see them.
Figure 10. Overall performance breakdown among participants with different levels of buckpassing: participants who chose to see AI decisions showed higher accuracy when the AI’s decisions were correct and lower accuracy when the AI’s decisions were incorrect, across all levels of buckpassing. When exposed to incorrect AI decisions, participants with high buckpassing scores (45%) showed lower accuracy than those with low buckpassing scores (49%). When exposed to correct AI decisions, participants with high buckpassing scores (84%) performed similarly to those with low buckpassing scores (85%). The n indicates the number of samples in each group and condition.